# Deep DeePC: Data-enabled predictive control with low or no online optimization using deep learning

### v1

@@ 2.2 Non-parametric representation of linear systems
Consider a discrete time-invariant linear (LTI) system in which the dynamic behaviors are described by the following state-space form:
@@
Let $L$, $T\in\mathbb{Z}_{>0}$ and $T\geq L$. Let $\mathbf{u}_{T}^{d}:=\{u^{d}\}_{1}^{T}\in\mathbb{R}^{n_{u}T}$ and $\mathbf{y}_{T}^{d}:=\{y^{d}\}_{1}^{T}\in\mathbb{R}^{n_{y}T}$ denote the input and output sequences for $T$ time instants, respectively. The superscript $d$ represents the sequences collected offline. For the input sequence $\mathbf{u}_{T}^{d}$, the Hankel matrix of depth $L$ is defined as follows:
@@
(Willems’ fundamental lemma [37]) Consider an LTI system (1) and assume this system is controllable. Let $\mathbf{u}_{T}^{d}:=\{u^{d}\}_{1}^{T}$ and $\mathbf{y}_{T}^{d}:=\{y^{d}\}_{1}^{T}$ be the $T$-step input and output sequences of system (1). Assume that the input sequence $\mathbf{u}_{T}^{d}$ is persistently exciting of order $L+n_{x}$. Then, any $L$-step sequences $\mathbf{u}_{L}:=\{u\}_{1}^{L}\in\mathbb{R}^{n_{u}L}$ and $\mathbf{y}_{L}:=\{y\}_{1}^{L}\in\mathbb{R}^{n_{y}L}$ are the input and output trajectories of system (1), if and only if
@@ 2.3 Data-enabled predictive control (DeePC)
where $U_{p}\in\mathbb{R}^{n_{u}T_{ini}\times(T-L+1)}$ denotes the past data which consist of the first $T_{ini}$ block rows of $\mathscr{H}_{L}(\mathbf{u}_{T}^{d})$; $U_{f}\in\mathbb{R}^{n_{u}N_{p}\times(T-L+1)}$ denotes the future data which consist of the last $N_{p}$ block rows of $\mathscr{H}_{L}(\mathbf{u}_{T}^{d})$ (similarly for $Y_{p}$ and $Y_{f}$).
@@ Acknowledgment
    This research is supported by the National Research Foundation, Singapore, and PUB, Singapore’s National Water Agency under its RIE2025 Urban Solutions and Sustainability (USS) (Water) Centre of Excellence (CoE) Programme, awarded to Nanyang Environment & Water Research Institute (NEWRI), Nanyang Technological University, Singapore (NTU). This research is also supported by Ministry of Education, Singapore, under its Academic Research Fund Tier 1 (RS15/21 & RG63/22). Any opinions, findings and conclusions or recommendations expressed in this material are those of the author(s) and do not reflect the views of the National Research Foundation, Singapore and PUB, Singapore’s National Water Agency.

### v2

@@ 2.2 Non-parametric representation of linear systems
Consider a discrete linear time-invariant (LTI) system in which the dynamic behaviors are described by the following state-space form:
@@
Let $L$, $T\in\mathbb{Z}_{>0}$ and $T\geq L$. Let $\mathbf{u}_{T}^{d}:=\{u^{d}\}_{1}^{T}\in\mathbb{R}^{n_{u}T}$ and $\mathbf{y}_{T}^{d}:=\{y^{d}\}_{1}^{T}\in\mathbb{R}^{n_{y}T}$ denote the input and output sequences for $T$ time instants, respectively, with the superscript $d$ indicating that the sequences are collected offline. For the input sequence $\mathbf{u}_{T}^{d}$, the Hankel matrix of depth $L$ is defined as follows:
@@
(Willems’ fundamental lemma [37]) Consider an LTI system (1) and assume this system is controllable. Consider that $\mathbf{u}_{T}^{d}:=\{u^{d}\}_{1}^{T}\in\mathbb{R}^{n_{u}T}$ and $\mathbf{y}_{T}^{d}:=\{y^{d}\}_{1}^{T}\in\mathbb{R}^{n_{y}T}$ are the T-step input and output sequences for system (1), respectively, and the input sequence $\mathbf{u}_{T}^{d}$ is persistently exciting of order $L+n_{x}$. Any $L$-step sequences $\mathbf{u}_{L}:=\{u\}_{1}^{L}\in\mathbb{R}^{n_{u}L}$ and $\mathbf{y}_{L}:=\{y\}_{1}^{L}\in\mathbb{R}^{n_{y}L}$ are the input and output trajectories of system (1), if and only if
@@ 2.3 Data-enabled predictive control (DeePC)
where $U_{p}\in\mathbb{R}^{n_{u}T_{ini}\times(T-L+1)}$ denotes the past data which consist of the first $n_{u}T_{ini}$ block rows of $\mathscr{H}_{L}(\mathbf{u}_{T}^{d})$; $U_{f}\in\mathbb{R}^{n_{u}N_{p}\times(T-L+1)}$ denotes the future data which consist of the last $n_{u}N_{p}$ block rows of $\mathscr{H}_{L}(\mathbf{u}_{T}^{d})$ (similarly for $Y_{p}$ and $Y_{f}$).
@@ Acknowledgment
    This research is supported by the National Research Foundation, Singapore, and PUB, Singapore’s National Water Agency under its RIE2025 Urban Solutions and Sustainability (USS) (Water) Centre of Excellence (CoE) Programme, awarded to Nanyang Environment & Water Research Institute (NEWRI), Nanyang Technological University, Singapore (NTU). This research is supported by Ministry of Education, Singapore, under its Academic Research Fund Tier 1 (RS15/21 & RG63/22), and Nanyang Technological University, Singapore (Start-Up Grant). Any opinions, findings and conclusions or recommendations expressed in this material are those of the author(s) and do not reflect the views of the National Research Foundation, Singapore and PUB, Singapore’s National Water Agency.